# Collision-based Watermark for Detecting Backdoor Manipulation in Federated Learning

### v2

@@ Abstract
As AI-generated content increasingly underpins real-world applications, its accompanying security risks, including privacy leakage and copyright infringement, have become growing concerns. In this context, Federated Learning (FL) offers a promising foundation for enhancing trustworthiness by enabling privacy-preserving collaborative learning over proprietary data. However, its practical adoption is critically threatened by backdoor-based model manipulation, where a small number of malicious clients can compromise the system and induce harmful content generation and decision-making. Although various detection methods have been proposed to detect such manipulation, we reveal that they are either disrupted by non-i.i.d. data distributions and random client participation, or misled by out-of-distribution (OOD) prediction bias, both of which are unique challenges in FL scenarios. To address these issues, we introduce a novel proactive detection method dubbed Coward, inspired by our discovery of multi-backdoor collision effects, in which consecutively planted, distinct backdoors significantly suppress earlier ones. Correspondingly, we modify the federated global model by injecting a carefully designed backdoor-collided watermark, implemented via regulated dual-mapping learning on OOD data. This design not only enables an inverted detection paradigm compared to existing proactive methods, thereby naturally counteracting the adverse impact of OOD prediction bias, but also introduces a low-disruptive training intervention that inherently limits the strength of OOD bias, leading to significantly fewer misjudgments. Extensive experiments on benchmark datasets show that Coward achieves state-of-the-art detection performance, effectively alleviates OOD prediction bias, and remains robust against potential adaptive manipulations. The code for our method is available at https://github.com/still2009/cowardFL.
@@
Backdoor Detection, Backdoor Manipulation, Federated Learning, Trustworthy Machine Learning, AI Security.
@@ 1 Introduction
AI-generated content, including text, images, numerical predictions, and categorical decisions, has become ubiquitous and now serves as a core component of modern intelligent applications [cao2025survey]. However, growing risks of harmful content generation [shen2025hatebench, ba2024exposing], malicious manipulation [chu2025jailbreakradar, shensynthetic], data privacy abuse [carlini2021extracting, chen2025survey], and copyright infringement [reuters2023_openai_copyright, liu2025harnessing], highlight an urgent need to develop trustworthy AI content generation and decision systems. As a promising starting point, Federated Learning (FL) [mcmahan2017fedavg, fl_survey24_pami] provides a trustworthy collaborative learning paradigm that enables the privacy-preserved utilization of cross-organization domain data. By keeping raw data locally owned and processed, FL naturally strengthens data privacy and copyright protection, while also unlocking the value of proprietary, professional, and domain-specific data that are inaccessible to models trained solely on generic public knowledge. Together, these properties position FL as an increasingly attractive solution for building trustworthy foundation models [fl_fm_nips, fan2025_fl_foudationmodel], and have motivated its wide exploration across various domains such as healthcare, finance, and edge intelligence [fl_nature_medi24, fl_finance_icse24, fl_edge_survey24].
@@
Figure 1: Our backdoor-collided detection versus the existing backdoor-coexistent method. By developing a backdoor-collided watermark, our method enables an inverted detection paradigm that remains effective under OOD prediction bias, a key challenge that limits current proactive detection methods.
@@
However, the privacy-preserving design of federated learning is a double-edged sword. While keeping raw data local effectively protects sensitive information, it also restricts the server’s ability to inspect participant behavior during training. This restricted observability creates a structural vulnerability to model-level manipulation, where the integrity of the federated global model can be compromised by covertly manipulated local models, most prominently through backdoor attacks [ba_survey_yiming, chameleon]. In general, backdoor attacks manipulate models to behave normally on benign inputs while producing harmful outputs only when specific trigger conditions are met, making them more stealthy and difficult to perceive. Once the manipulated local models are aggregated together with benign ones, the backdoor is seamlessly propagated to the global model, thereby manipulating the entire federated system [bagdasaryan2020backdoor, tifs25_fl_atk2].
@@
To detect and exclude backdoor-manipulated models, passive detection methods have been proposed [rieger2022deepsight, nguyen2022flame], which assume that malicious models expose observable model-level deviations as outliers. They rely on similarity-based measurements over either model parameters or model outputs and identify anomalous client models either through direct thresholding or by further applying clustering to expose structurally isolated outliers. However, we reveal that these methods are vulnerable to clients’ non-i.i.d. data distribution and participation randomness, which are common settings in FL scenarios [fl_survey24_pami]. These factors lead to highly dynamic and divergent benign local models that may appear as outliers, significantly interfering with detection. We refer to this limitation as the non-i.i.d. blurring challenge, as analyzed in Section 3.2.
@@
Recently, the pioneering work BackdoorIndicator [li2024backdoorindicator] proposed the first federated proactive detection strategy, which explicitly intervenes in the training process to induce manipulated models to exhibit behaviors distinctive from benign ones. Specifically, BackdoorIndicator injects OOD patterns specifically designed to be easily retained by manipulated clients while tending to be forgotten by benign ones, such that stronger pattern retention in a local model indicates a manipulated client. By shifting the detection basis from model discrepancies to differences in client responses to intervention actions, proactive methods naturally avoid the sensitivity to non-i.i.d. data that often undermines traditional defenses. However, due to the privacy constraints of FL that prevent access to in-distribution client data, proactive training interventions at the server are restricted to perform via OOD data. As a result, detection outcomes become susceptible to the OOD prediction bias of deep neural networks (DNNs) [nguyen2015ood_bias, guo2017ood_bias_calib, hein2019relu_ood_bias_theory], causing benign local models to be easily misidentified as malicious clients. We provide a detailed analysis in Section 3.3 and refer to this limitation as the OOD bias challenge.
@@
To tackle above challenges, we propose a simple yet effective Collision-based OOD WAtermark for Robust federated backdoor Detecton method, termed Coward. Our method is inspired by the finding of multiple backdoor manipulation collision effect where a consecutively planted different backdoor would significantly collide the one previously planted. To this end, we detect model manipulation by evaluating whether the server-injected conflicting global watermark has been erased during local training. Specifically, we implant a conflicting OOD watermark, where OOD samples embedded with a specific watermark trigger are redirected from their original class to a designated target class. Concretely, the OOD watermark is injected by jointly learning: (1) a regulated base OOD mapping that specifies a fixed class-wise correspondence from OOD samples to the task label space; and (2) a targeted watermark mapping that enforces an association between all triggered OOD samples and the watermark target class. In general, our method alleviates the OOD bias challenge through three aspects: (1) Unlike prior approaches, our collision-based detection treats high inspection accuracy as the signal of benign behavior, making it naturally compatible with the OOD bias–induced high confidence predictions; (2) Our watermark-based detection inspects the presence of the trigger-to-target mapping rather than the OOD sample-to-target mapping, mitigating the direct impact of OOD prediction bias. and (3) our method enforces explicit class-wise label assignments to OOD samples to reduce random prediction fluctuations; We demonstrate (1) and (2) in Figure1, and verify (3) in Figure7.
In summary, our contributions are three-fold:
  • We revisit existing detection methods against FL backdoor manipulation and identified two key challenges limiting their effectiveness: non-i.i.d. blurring challenge and OOD prediction bias unique characterized under FL sceanrios.
@@
  • We find an intriguing phenomenon, dubbed the multi-backdoor manipulation collision effect, where a consecutively planted different backdoor would significantly suppress the previous implanted one. Based on this finding, we design a simple yet effective proactive backdoor manipulation detection method, $i.e.$, Coward, that effectively mitigates OOD bias and achieves more accurate detection.
@@ 2.1 Federated Learning
  • Partial Client Participation [mcmahan2017fedavg]: In each training round, only a subset of clients is selected due to limited communication resources and dynamic client availability.
2.2 Backdoor Manipulation in Federated Learning
@@
Backdoor attacks aim to implant hidden behaviors into neural networks, causing them to function normally on clean data but misclassify inputs containing attacker-specified triggers [ba_survey_yiming, gu2020badnets]. Data poisoning is the most common backdoor injection method, where a subset of training samples is embedded with triggers and mislabeled to a target class [gu2020badnets]. Stealth and persistence are key objectives in backdoor attacks, motivating the design of invisible [nguyen2021wanet, Gao_attack24], semantic [bagdasaryan2021blind, xu2024csba_semantic_atk], optimized [geiping2020witches, sun2024backdoor] and sample-specific [iad, li2021invisible, zhu2025towards] triggers. However, these approaches are limited in FL, because adversaries can only manipulate a limited subset of clients and must ensure their backdoored updates survive the aggregation process to impact the global model. Specifically, FL-specific backdoor manipulation can be classified into following two categories.
@@
Single-client Manipulation. The first backdoor manipulations method for FL introduced the model replacement attack technic  [bagdasaryan2020backdoor], where a malicious client uploaded a scaled local update to dominate the aggregated model. However, the abnormally large update made it easily detectable under detection [zitengsun01pgd]. To improve stealth and persistence, follow-up works introduced more subtle manipulation strategies. For example, PGD [zitengsun01pgd] constrained malicious updates within an $\ell_{2}$-ball to evade norm-based detection, while Neurotoxin [Zhengming01Neurotoxin] injected backdoors into parameters associated with inactive neurons to minimize noticeable changes. Other works aimed to further blend backdoor patterns into normal behavior. For example, Edge-case manipulation [wang2020edgecase] utilized naturally occurring features as triggers, reducing the likelihood of detection, while Chameleon [chameleon] aligned backdoor and benign features using contrastive learning.
@@
Multi-client Manipulation. The DBA framework [xie2019dba] leveraged the distributed nature of FL to spread the backdoor across multiple clients, with each client using a distinct local trigger. These independently subtle triggers collectively formed an effective global backdoor, reducing the detectability of any individual client. Expanding on this idea, Gong et al. [gong2022coordinated_dba] proposed a coordinated multi-client manipulation method, in which each local trigger was optimized in a model-aware fashion. This coordination allowed the attackers to craft triggers that were both less suspicious and more effective, further amplifying the threat posed by distributed poisoning. In contrast, Non-Cooperative Backdoor Attacks (NBA) [nguyen2024nba] involved independent attackers using distinct triggers for different targets, yet still exhibited enhanced stealth comparable to coordinated methods.
2.3 Backdoor Detection in Federated Learning
@@
Backdoor Client Detection methods aim to identify and exclude manipulated local models during aggregation by quantifying the differences between benign and malicious clients, typically through distance-based metrics on model parameters. Clients whose update patterns deviate significantly from the majority are flagged as malicious. For instance, Multi-Krum [multiKrum] leveraged Euclidean distance, while Foolsgold [foolsgold], FLTrust [cao2021fltrust], and ShieldFL [ma2022shieldfl] adopted cosine similarity, all applied directly on full model weights. Rflbat [wang2022rflbat] extended this by computing distances in a top-$k$ PCA-projected space, aiming to reduce noise and improve efficiency, though at the cost of some information loss. Beyond model parameters, DeepSight [rieger2022deepsight] incorporated prediction-level discrepancies of local models, enabling a more comprehensive analysis. Extending beyond pure detection, Flame [nguyen2022flame] combined clustering-based client identification with gradient norm-based mitigation, positioning itself at the intersection of detection and mitigation strategies and demonstrating strong empirical performance. FLGuardian [tifs25_fl_defens_poison], a recent study addressing untagged backdoor manipulation, performed layer-wise model analysis using similarity metrics and clustering algorithms, reflecting a growing shift toward more fine-grained detection in the parameter space. However, these methods assume malicious models are outliers, making them sensitive to data heterogeneity and ineffective against stealthy attacks. Since these detection methods operate without any server-side intervention, we categorize them as passive detection. In contrast, proactive detection explicitly intervenes in FL training to facilitate backdoor identification. The recent pioneering work Indicator [li2024backdoorindicator] adopts this strategy by injecting OOD mappings into the global model, achieving improved robustness under data heterogeneity. However, it still suffers from high false-positive rates due to OOD prediction bias, leaving the design of effective proactive federated backdoor detection methods an open challenge.
3 Revisiting Existing Detection Methods and Multi-Backdoor Manipulation Interplay
In this section, we first introduce the threat model and revisit the limitations of existing detection methods. We then examine the interplay effects of multi-backdoor manipulation, which directly motivate our detection method.
@@ 3.2 Revisiting Passive Detection
Settings. We revisit gradient norm distribution for each client, under a highly non-i.i.d. partitioned CIFAR-10 dataset with label distribution generated by Dirichlet prior with $\alpha=0.3$, and 10 out of 100 random clients participation. One malicious client performs backdoor manipulation using a pixel-style trigger targeting class 0.
@@
Results. As shown in Figure 2(a), the suspicion toward manipulation is significantly reduced due to the severity of data heterogeneity: (1) in the first column, we present the size and label skewness of local datasets. The blue line indicates the dataset size, while the bar height represents the skewness, measured by the standard deviation of class-wise sample counts. We can observe that client data distributions are highly diverse. Notably, malicious clients do not exhibit abnormal characteristics in either size or skewness; their distributions fall within a typical range. In contrast, some benign clients show extreme distributions with both large data volumes and high skewness. (2) In the second column, we plot the mean and standard deviation of gradient norms uploaded by all 100 clients across training rounds, sorted by their mean values. The results reveal that the malicious client’ updates have slightly smaller gradient norms compared to most benign clients and appear inconspicuous amid the naturally diverse benign client updates. (3) In the third column, we visualize the top-2 PCA component of local and global models (based on the last-layer weights) during training. The results show that the manipulated models’ update direction and distance lie in a moderate region, appearing unharmful due to the diverse update directions introduced by the non-i.i.d. distribution. Figure 2(b) further presents a correlation analysis, revealing a clear association between model update norms and misjudgement of manipulations. Specifically, the x-axis shows the mean gradient norm, the y-axis indicates the number of misjudgments ($i.e.$, times a benign client was incorrectly flagged), and the bubble size reflects the standard deviation of the gradient norm. The evident upward trend suggests that clients with larger and more variable gradient norms are more susceptible to be misjudged. In conclusion, we have:

### v3

# Coward: Collision-based Watermark for Proactive Federated Backdoor Detection

@@ Abstract
Backdoor detection is currently the mainstream defense against backdoor attacks in federated learning (FL), where a small number of malicious clients can upload poisoned updates to compromise the federated global model. Existing backdoor detection techniques fall into two categories, passive and proactive, depending on whether the server proactively intervenes in the training process. However, both of them have inherent limitations in practice: passive detection methods are disrupted by common non-i.i.d. data distributions and random participation of FL clients, whereas current proactive detection methods are misled by an inevitable out-of-distribution (OOD) bias because they rely on backdoor coexistence effects. To address these issues, we introduce a novel proactive detection method dubbed Coward, inspired by our discovery of multi-backdoor collision effects, in which consecutively planted, distinct backdoors significantly suppress earlier ones. Correspondingly, we modify the federated global model by injecting a carefully designed backdoor-collided watermark, implemented via regulated dual-mapping learning on OOD data. This design not only enables an inverted detection paradigm compared to existing proactive methods, thereby naturally counteracting the adverse impact of OOD prediction bias, but also introduces a low-disruptive training intervention that inherently limits the strength of OOD bias, leading to significantly fewer misjudgments. Extensive experiments on benchmark datasets show that Coward achieves state-of-the-art detection performance, effectively alleviates OOD prediction bias, and remains robust against potential adaptive attacks. The code for our method is available at https://github.com/still2009/cowardFL.
@@
Backdoor Defense, Backdoor Attack, Federated Learning, AI Security, Trustworthy ML.
@@ 1 Introduction
Federated Learning (FL) [mcmahan2017fedavg, fl_big_survey, fl_survey24_pami] has emerged as a powerful paradigm for privacy-preserving machine learning, enabling collaborative model training across distributed clients without sharing raw data. By keeping data locally owned and processed, FL inherently strengthens privacy and copyright protection, and unlocks the value of proprietary, professional, domain-specific and personal data held by private organizations or individuals. This privacy-native property makes FL particularly suitable for securely building personalized and domain-specialized intelligence that goes beyond models trained solely on generic public data [tdsc25_fl_personalize, fl_fm_nips, fan2025_fl_foudationmodel]. As a result, FL has been extensively studied across a wide range of application domains, including healthcare, finance, and edge intelligence [fl_nature_medi24, fl_finance_icse24, fl_edge_survey24].
@@
Figure 1: Our backdoor-collided detection versus the existing backdoor-coexistent method. By developing a backdoor-collided watermark, our method enables an inverted detection paradigm that effectively mitigates misleading effects caused by the OOD prediction bias, a key challenge that limits current proactive detection methods.
@@
However, the privacy-preserving design of federated learning is a double-edged sword. While keeping raw data local effectively protects sensitive information, it also restricts the server’s ability to inspect participant behavior during training. This restricted observability creates an inevitable vulnerability to model-level attacks, where the integrity of the federated global model can be compromised by covertly manipulated local models, most prominently through backdoor attacks [ba_survey_yiming, chameleon, yang2024_ba_atk]. In general, backdoor attacks cause models to behave normally on benign inputs while producing harmful outputs only under specific trigger conditions, rendering the malicious behavior stealthy and difficult to detect. Once the malicious backdoor models are aggregated together with benign ones, the backdoor is seamlessly propagated to the global model, thereby manipulating the entire federated system [bagdasaryan2020backdoor, tdsc25_fba_atk1_durable, tdsc25_fba_atk2_collusive].
@@
To detect and exclude malicious backdoor models, passive detection methods have been proposed [rieger2022deepsight, nguyen2022flame], which do not intervene in the FL process and assume that malicious models manifest observable model-level deviations as outliers. They rely on similarity-based measurements over either model parameters or outputs and identify anomalous models either through direct thresholding or by further applying clustering to expose structurally isolated outliers. However, we reveal that these methods are vulnerable to client non-i.i.d. data distributions and participation randomness, which are common settings in FL scenarios [fl_survey24_pami]. These factors lead to highly dynamic and divergent benign local models that may appear as outliers, significantly interfering with detection. We refer to this limitation as the non-i.i.d. divergence challenge, as analyzed in Section 3.2.
@@
Recently, the pioneering work BackdoorIndicator [li2024backdoorindicator] proposed the first federated proactive detection strategy, which explicitly intervenes in the training process to induce malicious models to exhibit behaviors distinct from benign ones. Specifically, BackdoorIndicator injects OOD patterns designed to be easily retained by malicious client models while tending to be forgotten by benign ones, such that stronger pattern retention in a local model indicates a malicious client. By shifting the detection basis from model discrepancies to differences in client responses to intervention actions, proactive methods naturally avoid the sensitivity to non-i.i.d. distributions that often undermines passive detection methods. However, due to the privacy constraints of FL that prevent access to in-distribution client data, proactive training interventions at the server are restricted to operating via OOD data. As a result, detection outcomes become susceptible to the OOD prediction bias of deep neural networks (DNNs) [nguyen2015ood_bias, guo2017ood_bias_calib, hein2019relu_ood_bias_theory]. This bias causes benign local models to be easily misidentified as malicious clients. We provide a detailed analysis in Section 3.3 and refer to this limitation as the OOD bias challenge.
@@
To tackle above challenges, we propose a simple yet effective Collision-based OOD WAtermark for Robust federated backdoor Detection method, termed Coward. Our method is inspired by the finding of multi-backdoor collision effect where a consecutively planted distinct backdoor will significantly interfere with the one previously planted. To this end, we detect backdoor models by evaluating whether the server-injected conflicting global watermark has been erased during local training. Specifically, we implant a conflicting OOD watermark, where OOD samples embedded with a specific watermark trigger are redirected from their original class to a designated target class. Concretely, the OOD watermark is injected by jointly learning: (1) a regulated base OOD mapping that specifies a fixed class-wise correspondence from OOD samples to the task label space; and (2) a targeted watermark mapping that enforces an association between all triggered OOD samples and the watermark target class. In general, our method alleviates the OOD bias challenge through three aspects: (1) Unlike prior approaches, our collision-based detection treats high inspection accuracy as the signal of benign behavior, making it naturally compatible with the OOD bias–induced high confidence predictions; (2) Our watermark-based detection inspects the presence of the trigger-to-target mapping rather than the OOD sample-to-target mapping, mitigating the direct impact of OOD prediction bias. (3) Our method enforces explicit class-wise label assignments to OOD samples to reduce random prediction fluctuations. We demonstrate (1) and (2) in Figure 1, and verify (3) in Figure 7.
In summary, our contributions are threefold:
  • We revisit existing detection methods for FL backdoor attacks and identify two key challenges inherent to FL settings: non-i.i.d. divergence and OOD prediction bias. We show how these challenges undermine existing detection methods, leading to high benign-client misclassification.
@@
  • We find an intriguing phenomenon, dubbed the multi-backdoor collision effect, where a consecutively planted distinct backdoor can significantly suppress the previously implanted one. Based on this finding, we design a simple yet effective proactive backdoor detection method, $i.e.$, Coward, that effectively mitigates OOD bias and achieves more accurate detection.
@@ 2.1 Federated Learning
  • Partial Client Participation [mcmahan2017fedavg]: In each training round, only a subset of clients is randomly selected due to limited communication resources and dynamic client availability.
2.2 Backdoor Attacks in Federated Learning
@@
Backdoor attacks aim to implant hidden behaviors into neural networks, causing them to function normally on clean data but misclassify inputs containing attacker-specified triggers [ba_survey_yiming, gu2020badnets]. Data poisoning is the most common backdoor injection method, where a subset of training samples is embedded with triggers and mislabeled to a target class [gu2020badnets]. Stealth and persistence are key objectives in backdoor attacks, motivating the design of invisible [nguyen2021wanet, Gao_attack24], semantic [bagdasaryan2021blind, xu2024csba_semantic_atk], optimized [geiping2020witches, sun2024backdoor] and sample-specific [iad, li2021invisible, zhu2025towards] triggers. However, these approaches are limited in FL, because adversaries can only manipulate a limited subset of clients and must ensure their backdoored updates survive the aggregation process to impact the global model. Specifically, FL-specific backdoor attacks can be classified into following two categories.
@@
Single-client Attack. The first backdoor attack method for FL introduced the model replacement attack technic  [bagdasaryan2020backdoor], where a malicious client uploaded a scaled local update to dominate the aggregated model. However, the abnormally large update made it easily detectable under detection [zitengsun01pgd]. To improve stealth and persistence, follow-up works introduced more subtle attack strategies. For example, PGD [zitengsun01pgd] constrained malicious updates within an $\ell_{2}$-ball to evade norm-based detection, while Neurotoxin [Zhengming01Neurotoxin] injected backdoors into parameters associated with inactive neurons to minimize noticeable changes. Other works aimed to further blend backdoor patterns into normal behavior. For example, Edge-case attack [wang2020edgecase] utilized naturally occurring features as triggers, reducing the likelihood of detection, while Chameleon [chameleon] aligned backdoor and benign features using contrastive learning. Beyond stealthiness, a recent work [tdsc25_fba_atk1_durable] analyzes vulnerable layers in NLP models to design a more durable backdoor attack, achieving promising results on LSTM and GPT-2, while remaining limited to specific model architectures and task settings, yet offering valuable insights for general design for durability.
@@
Multi-client Attack. The DBA framework [xie2019dba] leveraged the distributed nature of FL to spread the backdoor across multiple clients, with each client using a distinct local trigger. These independently subtle triggers collectively formed an effective global backdoor, reducing the detectability of any individual client. Expanding on this idea, collusive multi-client attacks [gong2022coordinated_dba, tdsc25_fba_atk2_collusive] have been proposed, in which backdoor triggers are further optimized in a model-aware manner. Such learning-based coordination enables attackers to craft triggers that are both less suspicious and more effective, thereby amplifying the threat posed by distributed poisoning. In contrast, Non-Cooperative Backdoor Attacks (NBA) [nguyen2024nba] involved independent attackers using distinct triggers for different targets, yet still exhibited enhanced stealth comparable to coordinated methods.
2.3 Backdoor Defenses in Federated Learning
@@
Backdoor Client Detection methods aim to identify and exclude malicious local models during aggregation, typically by quantifying distance between benign and malicious client models to find outliers. For instance, Multi-Krum [multiKrum] leveraged Euclidean distance, while Foolsgold [foolsgold], FLTrust [cao2021fltrust], and ShieldFL [ma2022shieldfl] adopted cosine similarity, all applied directly on full model weights. Rflbat [wang2022rflbat] extended this by computing distances in a top-$k$ PCA-projected space, aiming to reduce noise and improve efficiency, though at the cost of some information loss. Beyond model parameters, DeepSight [rieger2022deepsight] incorporated prediction-level discrepancies of local models, enabling a more comprehensive analysis. Extending beyond pure detection, Flame [nguyen2022flame] combined clustering-based client identification with gradient norm-based mitigation, positioning itself at the intersection of detection and mitigation strategies and demonstrating strong empirical performance. FLGuardian [tifs25_fl_defens_poison], a recent study addressing untagged backdoor attack, performed layer-wise model analysis using similarity metrics and clustering algorithms, reflecting a growing shift toward more fine-grained detection in the parameter space. However, these methods assume malicious models are outliers, making them sensitive to data heterogeneity and ineffective against stealthy attacks.
Since all above detection methods operate without any server-side intervention, we categorize them as passive detection. In contrast, proactive detection explicitly intervenes in FL training to facilitate backdoor identification. The recent pioneering work Indicator [li2024backdoorindicator] adopts this strategy by injecting OOD mappings into the global model, achieving improved robustness under data heterogeneity. However, it still suffers from high false-positive rates due to OOD prediction bias. Recent work [tdsc25_fba_dfs2_proactive_poi] has also applied the proactive detection paradigm to poisoning attacks by amplifying directional deviations between malicious and benign updates through broadcasting a scaled and perturbed, detection-oriented surrogate of aggregated gradients. However, this approach relies on the explicit accuracy-degradation objective of poisoning attacks, which is less stealthy than backdoors and easier to detect, and further depends on parameter-space gradient comparisons that are susceptible to non-i.i.d. bias. In summary, effective proactive backdoor detection in FL remains an open challenge.
3 Revisiting Existing Detection Methods and Multi-Backdoor Interplay
In this section, we first introduce the threat model and revisit the limitations of existing detection methods. We then examine the interplay effects of multi-backdoor attack, which directly motivate our detection method.
@@ 3.2 Revisiting Passive Detection
Settings. We revisit gradient norm distribution for each client, under a highly non-i.i.d. partitioned CIFAR-10 dataset with label distribution generated by Dirichlet prior with $\alpha=0.3$, and 10 out of 100 random clients participation. One malicious client performs backdoor attack using a pixel-style trigger targeting class 0.
@@
Results. As shown in Figure 2(a), the suspicion toward attack is significantly reduced due to the severity of data heterogeneity: (1) in the first column, we present the size and label skewness of local datasets. The blue line indicates the dataset size, while the bar height represents the skewness, measured by the standard deviation of class-wise sample counts. We can observe that client data distributions are highly diverse. Notably, malicious clients do not exhibit abnormal characteristics in either size or skewness; their distributions fall within a typical range. In contrast, some benign clients show extreme distributions with both large data volumes and high skewness. (2) In the second column, we plot the mean and standard deviation of gradient norms uploaded by all 100 clients across training rounds, sorted by their mean values. The results reveal that the malicious client’ updates have slightly smaller gradient norms compared to most benign clients and appear inconspicuous amid the naturally diverse benign client updates. (3) In the third column, we visualize the top-2 PCA component of local and global models (based on the last-layer weights) during training. The results show that the backdoor models’ update direction and distance lie in a moderate region, appearing unharmful due to the diverse update directions introduced by the non-i.i.d. distribution. Figure 2(b) further presents a correlation analysis, revealing a clear association between model update norms and misjudgement of malicious clients. Specifically, the x-axis shows the mean gradient norm, the y-axis indicates the number of misjudgments ($i.e.$, times a benign client was incorrectly flagged), and the bubble size reflects the standard deviation of the gradient norm. The evident upward trend suggests that clients with larger and more variable gradient norms are more susceptible to be misjudged. In conclusion, we have: